Reduce the equation to a set of ODDE. using separation of variable


(del)²A+[k²+f(p)+1/p² g(B)+h(z)]A =0


1
Expert's answer
2021-05-17T16:55:00-0400

"\displaystyle (del)²A+[k²+f(p)+1/p² g(B)+h(z)]A =0\\
\textsf{The auxiliary equation is}\\
m^2 = -\left(k²+f(p)+1/p² g(B)+h(z)\right)\\

m = \pm i\sqrt{k²+f(p)+1/p² g(B)+h(z)}\\

\therefore A = C_1\cos\left(\sqrt{k²+f(p)+1/p² g(B)+h(z)}\right) + C_2\cos\left(\sqrt{k²+f(p)+1/p² g(B)+h(z)}\right)"